# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 2.1 Description of the algebra
We check inclusion first. Note that the condition $f_{-i}(v^{-1})=(-1)^{i}f_{i}(v)$ means that $\mathcal{A}$ is stable under the involution that sends $v$ to $v^{-1}$ and $w$ to $-w^{-1}$. Indeed, in terms of $u_{\pm}$ this involution interchanges $u_{+}$ and $u_{-}$. We check the second and third conditions by induction on the number of $H_{a}$ used in the multiplication. The base case is zero and one multiplications: $p(v)$ and $H_{a}=q^{\frac{a}{2}}(v^{a}u_{+}+v^{-a}u_{-})=q^{\frac{-a}{2}}(u_{+}v^{a}+u_{-}v^{-a})$. The polynomial $p(v)$ has no poles. For $H_{a}$ we write $u_{+}v^{a}+u_{-}v^{-a}=wv^{a}(v-v^{-1})^{-1}+w^{-1}v^{-a}(v-v^{-1})^{-1}$, so that $f_{1}=\frac{v^{a}}{v-v^{-1}}$ and $f_{-1}=\frac{v^{-a}}{v-v^{-1}}$. All $f_{i}^{(j)}$ are zero except $f_{1}^{(0)}=\frac{1}{2}$, $f_{1}^{(0,-)}=\frac{(-1)^{a}}{2}$, $f_{-1}^{(0)}=\frac{1}{2}$, $f_{1}^{(0,-)}=\frac{(-1)^{-a}}{2}$. It follows that both the second and third condition are satisfied.
@@
We turn to the induction step. We will prove that if $x=\sum w^{i}f_{i}$ satisfies the conditions then $H_{a}x=(wg_{1}+w^{-1}g_{-1})x$ also satisfies the second and the third condition. Here $g_{\pm 1}$ are rational functions in $v$ with at most simple pole at $v=\pm 1$ satisfying $g_{1}^{(1,\pm)}=g_{-1}^{(1,\pm)}$. We also have $g_{-1}=-g_{1}(v^{-1})$.
@@
$H_{a}x=(wg_{1}+w^{-1}g_{-1})(\sum w^{i}f_{i})=\sum w^{i}(g_{1}(q^{1-i}v)f_{i-1}+g_{-1}(q^{-1-i}v)f_{i+1})=\sum w^{i}h_{i}.$ (2.1)
Note that $g_{\pm 1}(q^{-i}v)$ has at most simple poles at ±$\pm q^{i}$ with residue $q^{i}g_{\pm 1}^{(0,\pm)}$. Using the fact that $f_{i}$ is regular at $q^{i}$ we see that $g_{\pm 1}(q^{-i}v)f_{i}$ has simple poles.
Hence $h_{i}$ has simple poles. It remains to prove that for all $i$ and $k\geq 0$ we have $h_{i+k}^{(i,\pm)}=(-1)^{k+1}h_{i-k}^{(i,\pm)}$. We will prove it for ±=$\pm$=+± = +, and for ±$\pm$=-± = - the proof is the same.
For $j\neq i\pm 1$ the number $h_{i}^{(j)}$ can be computed as
$h_{i}^{(j)}=g_{1}(q^{j+1-i})f_{i-1}^{(j)}+g_{-1}(q^{j-1-i})f_{i+1}^{(j)}.$ (2.2)
Hence for $k\neq 1$ we have
$h_{i+k}^{(i)}=g_{1}(q^{k+1})f_{i+k-1}^{(i)}+g_{-1}(q^{k-1})f_{i+k+1}^{(i)}=-g_{-1}(q^{-k-1})(-1)^{k}f_{i-k+1}^{(i)}-g_{1}(q^{1-k})(-1)^{k}f_{i-k-1}^{(i)}=(-1)^{k+1}h_{i-k}^{(i)}.$
For $j=i+1$ we have
$h_{i}^{(i+1)}=g_{1}(q^{2})f_{i-1}^{(i+1)}+q^{i+1}g_{-1}^{(0)}f_{i+1}(q^{i+1}),$
and for $j=i-1$ we have
$h_{i}^{(i-1)}=q^{i-1}g_{1}^{(0)}f_{i-1}(q^{i-1})+g_{-1}(q^{-2})f_{i+1}^{(i-1)}$ (2.3)
. Using this we check the condition for $k=1$:
$h_{i+1}^{(i)}=q^{i}g_{1}^{(0)}f_{i}(q^{i})+g_{-1}(q^{-2})f_{i+2}^{(i)}=q^{i}g_{-1}^{(0)}f_{i}(q^{i})+(-g_{1}(q^{2}))\cdot(-f_{i-2}^{(i)})=h_{i-1}^{(i)}.$
This finishes the proof of inclusion. Now we prove that any element $a$ of $W$ satisfying the conditions of the proposition statement belongs to $\mathcal{A}$.
First we prove that any element $a=\sum_{i=-N}^{N}w^{i}f_{i}$ such that $f_{i}$ have no poles and satisfy $f_{-i}(v^{-1})=(-1)^{i}f_{i}(v)$ belongs to $\mathcal{A}$. As noted above, $S_{b}=q^{\tfrac{b}{2}}H_{b}=w\frac{v^{b}}{v-v^{-1}}+w^{-1}\frac{v^{-b}}{v-v^{-1}}$. We have $T_{b}=S_{b+1}-S_{b-1}=wv^{b}-w^{-1}v^{-b}$. It follows that for any $l\geq 1$ the element $T_{0}^{l-1}T_{b}$ can be written as $w^{l}v^{b}+(-1)^{l}w^{-l}v^{-b}+\sum_{i=1-l}^{l-1}w^{i}f_{i}$. Now we can prove that $a$ belongs to $\mathcal{A}$ using induction on $N$.
[…]
@@
Hence the number $f_{N}^{(j,\pm)}$ can be nonzero only if $0\leq j<N$. We will use equation (2.1) for the element $x=\sum_{i=1-N}^{N-1}w^{i}f_{i}$ and apply the induction hypothesis for $x$ and $N-1$. This means that for any sequence of numbers $c_{0}^{\pm},\ldots,c_{N-2}^{\pm}$ we can find $x\in\mathcal{A}$ such that $f_{N-1}^{(j,\pm)}=c_{j}^{\pm}$. We want to prove the same for $h_{N}^{(0,\pm)},\ldots,h_{N}^{(N-1,\pm)}$. As above, we assume that ±=$\pm$=+± = +, and the choices for ±$\pm$=-± = - are done independently.
Setting $i=N$, $j<N-1$ in equation (2.2) we get
$h_{N}^{(j)}=g_{1}(q^{j+1-N})f_{N-1}^{(j)}=g_{1}(q^{j+1-N})c_{j}.$
Setting $i=N$ in equation (2.3) we get
$h_{N}^{(N-1)}=q^{N-1}g_{1}^{(0)}f_{N-1}(q^{N-1}).$
Using induction hypothesis again we get that if we fix $c_{0},\ldots,c_{N-2}$, the number $f_{N-1}(q^{N-1})$ can be anything. Since $g_{1}^{(0)}$ is nonzero, we get that $h_{N}^{(0)},\ldots,h_{N}^{(N-1)}$ can be any sequence of numbers. This finishes the proof of the statement. ∎
2.2 Twisted traces
For the next proposition, note the following. Let $a=\sum w^{i}f_{i}$ be an element of $\mathcal{A}$. Using 2.1.1 we get $f_{0}(v)=f_{0}(v^{-1})$, $f_{0}^{(0,\pm)}=0$, and there are no other restrictions on the element $f_{0}$. Hence we have a surjective linear map $a\mapsto f_{0}$ from $\mathcal{A}$ to
$A_{0}=\mathbb{C}[v+v^{-1}]\oplus\sum_{k\neq 0}\frac{1}{(v-q^{k})(v^{-1}-q^{k})}\oplus\sum_{k\neq 0}\frac{1}{(v+q^{k})(v^{-1}+q^{k})}.$
Proposition 2.2.1.
@@ 2.2 Twisted traces
  • For an element $a=\sum_{i}w^{2i+1}f_{2i+1}(v)$ the value $T(a)$ is uniquely defined by $T(w-w^{-1})$ and $T(wv-w^{-1}v^{-1})$. If $a=wf_{1}+w^{-1}f_{-1}$ then $T(a)=C_{+}f_{1}(\sqrt{q})+C_{-}f_{1}(-\sqrt{q})$ for some numbers $C_{+},C_{-}$.
  • There exists a linear map $S\colon A_{0}\to\mathbb{C}$ such for any $a\in\mathcal{A}$ with $a=\sum_{i}w^{2i}f_{2i}$ we have $T(a)=S(f_{0})$.
@@
First we note that $g$ is restricted from the following automorphism of $W$, which we will also denote by $g$: $g(v)=v$, $g(w)=q^{\frac{k}{2}}v^{k}w=q^{-\frac{k}{2}}wv^{k}$.
We check the trace condition for a symmetric Laurent polynomial $h(v)$, it is preserved by $g$. Then for any $a=\sum_{i=-N}^{N}w^{i}f_{i}\in\mathcal{A}$ we have
$T(h(v)a)=T(ah(v)).$
We have $ah(v)=\sum w^{i}f_{i}h$ and $h(v)a=\sum h(v)w^{i}f_{i}=\sum w^{i}h(q^{-i}v)f_{i}$. It follows that for any $a\in\mathcal{A}$ and symmetric Laurent polynomial $g$ we have
$T\big{(}\sum w^{i}f_{i}(h(v)-h(q^{-i}v))\big{)}=0.$
@@
It follows from Proposition 2.1.1 that for any rational function $f_{N}$ with simple poles at ±$\pm 1,\ldots,\pm q^{N-1}$ we can find $f_{N-2},\ldots,f_{-N}$ such that $a=\sum w^{i}f_{i}$ belongs to $\mathcal{A}$. Note that $h(v)-h(q^{-N}v)$ is zero at $v=\pm q^{\frac{N}{2}}$ and this is the only condition: for $h=v+v^{-1}$ we have $h(v)-h(q^{-N}v)=(1-q^{-N})v+(1-q^{N})v^{-1}=(1-q^{-N})(v-q^{N}v^{-1})$. Hence for even $N$ the product $f_{N}(h(v)-h(q^{-N}v))$ can be any rational function with simple poles at ±$\pm 1,\ldots,\widehat{\pm q^{\frac{N}{2}}},\ldots,\pm q^{N-1}$ and for odd $N$ the product $f_{N}(h(v)-h(q^{N}v))$ can be any rational function with simple poles at ±$\pm 1,\ldots,\pm q^{N-1}$ with zero at ±$\pm q^{\frac{N}{2}}$.
By induction it follows that for an element $a=\sum_{i}w^{2i+1}f_{2i+1}(v)$ we have
$T(a)=\sum_{i\geq 0,\varepsilon=\pm}c_{2i+1,\varepsilon}f_{2i+1}(\varepsilon q^{\frac{2i+1}{2}}).$
Now we use the trace condition for $a=w-w^{-1}$ and $b=w^{2M}+w^{-2M}$ or $w^{2M}v+w^{-2M}v^{-1}$. We have $g(a)=q^{-\frac{m}{2}}(wv^{m}-w^{-1}v^{-m})$. Hence
@@
$bg(a)=$ $q^{-\frac{m}{2}}(w^{2M+1}v^{m}-w^{2M-1}v^{-m}+w^{1-2M}v^{m}-w^{-2M-1}v^{-m}),$
and a similar equation for $b=w^{2M}v+w^{-2M}v^{-1}$. Using $T(ab)=T(bg(a))$ we can express $c_{2M+1,\pm}$ using $c_{2M-1,\pm}$. Hence for an element $a=\sum_{i}w^{2i+1}f_{2i+1}(v)$ the value $T(a)$ is uniquely defined by $T(w-w^{-1})$ and $T(wv-w^{-1}v^{-1})$.
Let $N=2M$ be even. We have just showed that for any $x=\sum_{i=-M}^{M}w^{2i}p_{2i}(v)\in\mathcal{A}$ such that $p_{N}$ has no poles at ±$\pm q^{M}$, there exists $y=\sum_{i=1-N}^{N-1}w^{i}q_{i}(v)$ such that $T(x)=T(y)$ for all $g$-twisted traces $T$. Moreover, $p_{0}=h_{0}$.
Let $a=\sum_{i=-M}^{M}w^{2i}p_{2i}(v)\in\mathcal{A}$ be any element. By induction, it follows that $T(a)$ is uniquely defined by the values of $T$ on $\mathbb{C}[v+v^{-1}]$ and the values of $T$ on
@@
We see that the constant terms of these element form a basis of $A_{0}$. Define $S(f(v))=T(f(v))$ for $f\in\mathbb{C}[v+v^{-1}]$ and
$S\big{(}\frac{1}{v\pm q^{k}}+\frac{1}{v^{-1}\pm q^{k}}\big{)}=T\big{(}w^{2k}\frac{1}{v\pm q^{k}}+\frac{(-1)^{k+1}}{v\pm q^{k}}+\frac{(-1)^{k+1}}{v^{-1}\pm q^{k}}+w^{-2k}\frac{1}{v^{-1}\pm q^{k}}\big{)}.$
The proposition follows.
Proposition 2.2.2.
Let $S,T$ be as above. Then
$S(F(z))=\int_{\varepsilon S^{1}}F(z)\omega(z)\frac{dz}{z},$
where $S^{1}$ is the unit circle, $\omega(z)$ is a holomorphic function on $\mathbb{C}^{\times}$ such that $\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z)$ and $\omega(z)=\omega(z^{-1})$, and $\varepsilon$ is any number smaller than all of the absolute values of the poles of $F$. In particular, the dimension of space of traces is (around) $\frac{m}{2}$.
@@
We use the trace condition $T(ba)=T(ag(b))$ for $a=\sum w^{2i+1}f_{2i+1}$ and $b=w\frac{v^{k}}{v-v^{-1}}+w^{-1}\frac{v^{-k}}{v-v^{-1}}$. Since $ba$ and $ag(b)$ are symmetric with respect to $w\mapsto-w$, we can compute $T(ba)$ and $T(ag(b))$ using Proposition 2.2.1. We note that the only condition on $f_{1}(v)$ is that it is regular at $v=\pm q$. We have $f_{-1}(v)=-f_{1}(v^{-1})$. Denote $f_{1}(v)=F(v)$, so that $f_{-1}(v)=-F(v^{-1})$. We have
$[w^{0}](ba)=w\frac{v^{k}}{v-v^{-1}}\cdot(-w^{-1}F(v^{-1}))+w^{-1}\frac{v^{-k}}{v-v^{-1}}wF(v)=q^{k}\bigg{(}\frac{-v^{k}}{qv-q^{-1}v^{-1}}F(v^{-1})+\frac{v^{-k}}{q^{-1}v-qv^{-1}}F(v)\bigg{)}.$
Note that $g(b)=q^{-\frac{m}{2}}(w\frac{v^{k+m}}{v-v^{-1}}+w^{-1}\frac{v^{-k-m}}{v-v^{-1}})$. Hence
$[w^{0}](ag(b))=q^{-\frac{m}{2}}\bigg{(}wF(v)w^{-1}\frac{v^{-k-m}}{v-v^{-1}}-w^{-1}F(v^{-1})w\frac{v^{k+m}}{v-v^{-1}}\bigg{)}=q^{-\frac{m}{2}}\bigg{(}F(qv)\frac{v^{-k-m}}{v-v^{-1}}-F(qv^{-1})\frac{v^{k+m}}{v-v^{-1}}\bigg{)}.$

### v2

@@ 3. $(3)$
We should check that any element of $\mathcal{A}$ satisfies the conditions (1)–(3) and that any element satisfying conditions (1)–(3) belongs to $\mathcal{A}$.
Elements of $\mathcal{A}$ satisfy the conditions. Note that an element $a=\sum w^{i}f_{i}\in W$ satisfies the condition $f_{-i}\bigl(v^{-1}\bigr)=(-1)^{i}f_{i}(v)$ if and only if $a$ is fixed by the involution $\tau$ that sends $v$ to $v^{-1}$ and $w$ to $w^{-1}$. In terms of $u_{\pm}$, we have $\tau(u_{+})=u_{-}$ and $\tau(u_{-})=u_{+}$, so that $\tau(H_{a})=H_{a}$. It follows that all generators of $\mathcal{A}$ are fixed by $\tau$. This shows that the first condition is satisfied.
We check the second and third conditions by induction on the number of $H_{a}$ used in the multiplication. The base case is zero and one multiplications: $p(v)$ and
$H_{a}=q^{\frac{a}{2}}(v^{a}u_{+}+v^{-a}u_{-})=\smash{q^{\frac{-a}{2}}(u_{+}v^{a}+u_{-}v^{-a})}.$
The polynomial $p(v)$ has no poles. For $H_{a}$, we write
$u_{+}v^{a}+u_{-}v^{-a}=\smash{wv^{a}\bigl(v-v^{-1}\bigr)^{-1}+w^{-1}v^{-a}\bigl(v-v^{-1}\bigr)^{-1}},$
so that $f_{1}=\frac{v^{a}}{v-v^{-1}}$ and $f_{-1}=\frac{v^{-a}}{v-v^{-1}}$. All $f_{i}^{(j)}$ are zero except
$f_{1}^{(0)}=\tfrac{1}{2},\qquad f_{1}^{(0,-)}=\tfrac{(-1)^{a}}{2},\qquad f_{-1}^{(0)}=\tfrac{1}{2},\qquad f_{1}^{(0,-)}=\tfrac{(-1)^{-a}}{2}.$
It follows that both the second and third condition are satisfied.
@@
We turn to the induction step. We will prove that if $x=\sum w^{i}f_{i}$ satisfies the conditions then
$H_{a}x=(wg_{1}+w^{-1}g_{-1})x$
also satisfies the second and third conditions. Here $g_{\pm 1}=\frac{v^{\pm a}}{v-v^{-1}}$ are rational functions in $v$ with at most simple pole at $v=\pm 1$ satisfying $g_{1}^{(1,\pm)}=g_{-1}^{(1,\pm)}$. We also have $g_{-1}=-g_{1}\bigl(v^{-1}\bigr)$.
@@
$H_{a}x$ $=\bigl(wg_{1}+w^{-1}g_{-1}\bigr)\Big(\sum w^{i}f_{i}\Big)=\sum w^{i}\bigl(g_{1}\bigl(q^{1-i}v\bigr)f_{i-1}+g_{-1}\bigl(q^{-1-i}v\bigr)f_{i+1}\bigr)$
$=\sum w^{i}h_{i}.$ (7.1)
Note that $g_{\pm 1}\bigl(q^{-i}v\bigr)$ has at most simple poles at $\pm q^{i}$ with residue $q^{i}g_{\pm 1}^{(0,\pm)}$. Using the fact that $f_{i}$ is regular at $q^{i}$, we see that $g_{\pm 1}\bigl(q^{-i}v\bigr)f_{i}$ has simple poles.
Hence, $h_{i}$ has simple poles. It remains to prove that for all $i$ and $k\geq 0$, we have $\smash{h_{i+k}^{(i,\pm)}}=\smash{(-1)^{k+1}h_{i-k}^{(i,\pm)}}$. We will prove it for $\pm=+$, and for $\pm=$- the proof is the same.
For $j\neq i\pm 1$, the residue $h_{i}^{(j)}$ can be computed as
$h_{i}^{(j)}=g_{1}\bigl(q^{j+1-i}\bigr)f_{i-1}^{(j)}+g_{-1}\bigl(q^{j-1-i}\bigr)f_{i+1}^{(j)}.$ (7.2)
Hence, for $k\neq 1$, we have
$h_{i+k}^{(i)}$ $=g_{1}\bigl(q^{k+1}\bigr)f_{i+k-1}^{(i)}+g_{-1}\bigl(q^{k-1}\bigr)f_{i+k+1}^{(i)}$
$=-g_{-1}\bigl(q^{-k-1}\bigr)(-1)^{k}f_{i-k+1}^{(i)}-g_{1}\bigl(q^{1-k}\bigr)(-1)^{k}f_{i-k-1}^{(i)}=(-1)^{k+1}h_{i-k}^{(i)}.$
For $j=i+1$, we have
$h_{i}^{(i+1)}=g_{1}\bigl(q^{2}\bigr)f_{i-1}^{(i+1)}+q^{i+1}g_{-1}^{(0)}f_{i+1}\bigl(q^{i+1}\bigr),$
and for j=i1j=i-1 we have
$h_{i}^{(i-1)}=q^{i-1}g_{1}^{(0)}f_{i-1}\bigl(q^{i-1}\bigr)+g_{-1}\bigl(q^{-2}\bigr)f_{i+1}^{(i-1)}.$ (7.3)
Using this, we check the condition for $k=1$
$h_{i+1}^{(i)}=q^{i}g_{1}^{(0)}f_{i}\bigl(q^{i}\bigr)+g_{-1}\bigl(q^{-2}\bigr)f_{i+2}^{(i)}=q^{i}g_{-1}^{(0)}f_{i}\bigl(q^{i}\bigr)+\bigl(-g_{1}\bigl(q^{2}\bigr)\bigr)\cdot\bigl(-f_{i-2}^{(i)}\bigr)=h_{i-1}^{(i)}.$
This finishes the proof of inclusion.
Conditions (1)–(3) imply being inside $\mathcal{A}$. First we prove that any element $a=\sum_{i=-N}^{N}w^{i}f_{i}$ such that $f_{i}$ have no poles and satisfy $f_{-i}\bigl(v^{-1}\bigr)=(-1)^{i}f_{i}(v)$ belongs to $\mathcal{A}$. Consider the elements
$S_{b}=q^{\tfrac{b}{2}}H_{b}=w\frac{v^{b}}{v-v^{-1}}+w^{-1}\frac{v^{-b}}{v-v^{-1}}\qquad\text{and}\qquad T_{b}=S_{b+1}-S_{b-1}=wv^{b}-w^{-1}v^{-b}.$
Then for any $l\geq 1$, we have
$T_{0}^{l-1}T_{b}=w^{l}v^{b}+(-1)^{l}w^{-l}v^{-b}+\sum_{i=1-l}^{l-1}w^{i}f_{i}.$
Now we can prove that $a$ belongs to $\mathcal{A}$ using the induction on $N$.
Now let $a=\sum_{i=-N}^{N}w^{i}f_{i}$ be an element of $W_{0}$ such that $f_{i}$ has simple poles satisfying the conditions (1)–(3) above. We use induction on $N$, the base case $N=0$ being clear. We note that it is enough to find an element of $\mathcal{A}$ that cancels the poles of $f_{N}$, as then the poles of $f_{-N}$ will be also canceled by symmetry. We claim that $f_{N}^{(j,\pm)}$ is zero when $j>N$ or $j<0$. Indeed, for j>N{j>N} we use the third condition for i=j{i=j}, k=jNk=j-N to get
$f^{(j)}_{2j-N}=(-1)^{j-N+1}f^{(j)}_{N}$.
Since 2jN>N{2j-N>N}, we have $f_{2j-N}=0$, hence the left-hand side is zero. For $j<0$, we use the third condition for i=j{i=j}, k=Njk=N-j to get
$f_{N}^{(j)}=(-1)^{N-j+1}f_{2j-N}^{(j)}=0$
since 2jN<N2j-N<-N.
@@
Hence, the number $f_{N}^{(j,\pm)}$ can be nonzero only if $0\leq j<N$. We use equation (7.1) for the element $x=\sum_{i=1-N}^{N-1}w^{i}f_{i}$ and apply the induction hypothesis for $x$ and N1N-1. This means that for any sequence of numbers $c_{0}^{\pm},\ldots,c_{N-2}^{\pm}$, we can find $x\in\mathcal{A}$ such that
$f_{N-1}^{(j,\pm)}=c_{j}^{\pm}.$
We want to prove the same for $h_{N}^{(0,\pm)},\ldots,h_{N}^{(N-1,\pm)}$. As above, we assume that $\pm=+$, and the choices for $\pm=$- are done independently.
Setting $i=N$, j<N1j<N-1 in equation (7.2), we get
$h_{N}^{(j)}=g_{1}\bigl(q^{j+1-N}\bigr)f_{N-1}^{(j)}=g_{1}\bigl(q^{j+1-N}\bigr)c_{j}.$
Setting $i=N$ in equation (7.3) we get $h_{N}^{(N-1)}=q^{N-1}g_{1}^{(0)}f_{N-1}\bigl(q^{N-1}\bigr)$. Using induction hypothesis again, we get that if we fix $c_{0},\ldots,c_{N-2}$, the number $f_{N-1}\bigl(q^{N-1}\bigr)$ can be anything. Since $g_{1}^{(0)}$ is nonzero, we get that $h_{N}^{(0)},\ldots,h_{N}^{(N-1)}$ can be any sequence of numbers. This finishes the proof of the statement. ∎
8 Twisted traces
The algebra $\mathcal{A}=\mathcal{A}_{0}\oplus\mathcal{A}_{1}$ is $\mathbb{Z}/2$-graded: the even piece $\mathcal{A}_{0}$ consists of elements $a=\sum_{i\in 2\mathbb{Z}}w^{i}f_{i}$, the odd piece $\mathcal{A}_{1}$ consists of elements $a=\sum_{i\in 2\mathbb{Z}+1}w^{i}f_{i}$. We will use this grading below.
For the next proposition, note the following. Let $a=\sum w^{i}f_{i}$ be an element of $\mathcal{A}$. Using 7.6, we get $f_{0}(v)=f_{0}\bigl(v^{-1}\bigr)$, $f_{0}^{(0,\pm)}=0$, and there are no other restrictions on the element $f_{0}$. Hence, we have a surjective linear map $a\mapsto f_{0}$ from $\mathcal{A}$ to the vector space
$V=\mathbb{C}\big[v+v^{-1}\big]\oplus\bigoplus_{k\neq 0}\mathbb{C}\frac{1}{\bigl(v-q^{k}\bigr)\bigl(v^{-1}-q^{k}\bigr)}\oplus\bigoplus_{k\neq 0}\mathbb{C}\frac{1}{\bigl(v+q^{k}\bigr)\bigl(v^{-1}+q^{k}\bigr)}.$
Proposition 8.1.
@@ 8 Twisted traces
  • For an element $a\in\mathcal{A}_{1}$, the value $T(a)$ is uniquely defined by $T\bigl(w-w^{-1}\bigr)$ and $T\bigl(wv-w^{-1}v^{-1}\bigr)$. If $a=wf_{1}+w^{-1}f_{-1}$ then $T(a)=C_{+}f_{1}(\sqrt{q})+C_{-}f_{1}(-\sqrt{q})$ for some numbers $_{+}CC_{-}$.
  • There exists a linear map $S\colon V\to\mathbb{C}$ such for any $a\in\mathcal{A}_{0}$, we have $T(a)=S(f_{0})$.
@@
Recall that $g$ is restricted from the following automorphism of $W$, which we will also denote by $g$: $g(v)=v$, $g(w)=q^{\frac{k}{2}}v^{k}w=q^{-\frac{k}{2}}wv^{k}$.
We check the trace condition for a symmetric Laurent polynomial $h(v)$, it is preserved by $g$. Then for any $a=\sum_{i=-N}^{N}w^{i}f_{i}\in\mathcal{A}$, we have $T(h(v)a)=T(ah(v))$. We have $ah(v)=\sum w^{i}f_{i}h$ and $h(v)a=\sum h(v)w^{i}f_{i}=\sum w^{i}h\bigl(q^{-i}v\bigr)f_{i}$. It follows that for any $a\in\mathcal{A}$ and symmetric Laurent polynomial $g$, we have
$T\Big(\sum w^{i}f_{i}\bigl(h(v)-h\bigl(q^{-i}v\bigr)\bigr)\Big)=0.$
@@
It follows from Proposition 7.6 that for any rational function $f_{N}$ with simple poles at $\pm 1,\ldots,\pm q^{N-1}$ we can find $f_{N-2},\ldots,f_{-N}$ such that $a=\sum w^{i}f_{i}$ belongs to $\mathcal{A}$. Note that $h(v)-h\bigl(q^{-N}v\bigr)$ is zero at $v=\pm q^{\frac{N}{2}}$ and this is the only condition: for $h=v+v^{-1}$, we have
$h(v)-h\bigl(q^{-N}v\bigr)=\bigl(1-q^{-N}\bigr)v+\bigl(1-q^{N}\bigr)v^{-1}=\bigl(1-q^{-N}\bigr)\bigl(v-q^{N}v^{-1}\bigr).$
Hence, for even $N$, the product $f_{N}\bigl(h(v)-h\bigl(q^{-N}v\bigr)\bigr)$ can be any rational function with simple poles at
$\pm 1,\ldots,\widehat{\pm q^{\frac{N}{2}}},\ldots,\pm q^{N-1},$
and for odd $N$ the product $f_{N}\bigl(h(v)-h\bigl(q^{N}v\bigr)\bigr)$ can be any rational function vanishing at $\pm q^{\frac{N}{2}}$ with simple poles at $\pm 1,\ldots,\pm q^{N-1}$.
By induction, it follows that for an element $a=\sum_{i}w^{2i+1}f_{2i+1}(v)$ we have
$T(a)=\sum_{i\geq 0,\varepsilon=\pm}c_{2i+1,\varepsilon}f_{2i+1}\bigl(\varepsilon q^{\frac{2i+1}{2}}\bigr).$
Now we use the trace condition for $a=w-w^{-1}$ and $b=w^{2M}+w^{-2M}$ or $w^{2M}v+w^{-2M}v^{-1}$. We have $g(a)=q^{-\frac{m}{2}}\bigl(wv^{m}-w^{-1}v^{-m}\bigr)$. Hence,
@@
$bg(a)=q^{-\frac{m}{2}}\bigl(w^{2M+1}v^{m}-w^{2M-1}v^{-m}+w^{1-2M}v^{m}-w^{-2M-1}v^{-m}\bigr),$
and a similar equation for $b=w^{2M}v+w^{-2M}v^{-1}$. Using $T(ab)=T(bg(a))$, we can express $c_{2M+1,\pm}$ using $c_{2M-1,\pm}$. Hence, for an element $a=\sum_{i}w^{2i+1}f_{2i+1}(v)$, the value $T(a)$ is uniquely defined by $T\bigl(w-w^{-1}\bigr)$ and $T\bigl(wv-w^{-1}v^{-1}\bigr)$.
Let $N=2M$ be even. We have just showed that for any $x=\sum_{i=-M}^{M}w^{2i}p_{2i}(v)\in\mathcal{A}$ such that $p_{N}$ has no poles at $\pm q^{M}$, there exists $y=\sum_{i=1-N}^{N-1}w^{i}q_{i}(v)$ such that $T(x)=T(y)$ for all $g$-twisted traces $T$. Moreover, $p_{0}=h_{0}$.
Let $a=\sum_{i=-M}^{M}w^{2i}p_{2i}(v)\in\mathcal{A}$ be any element. By induction, it follows that $T(a)$ is uniquely defined by the values of $T$ on $\mathbb{C}\big[v+v^{-1}\big]$ and the values of $T$ on
@@
We see that the constant terms of these element form a basis of $V$. Define $S(f(v))=T(f(v))$ for $f\in\mathbb{C}\big[v+v^{-1}\big]$ and
$S\left(\frac{1}{v\pm q^{k}}+\frac{1}{v^{-1}\pm q^{k}}\right)=T\left(w^{2k}\frac{1}{v\pm q^{k}}+\frac{(-1)^{k+1}}{v\pm q^{k}}+\frac{(-1)^{k+1}}{v^{-1}\pm q^{k}}+w^{-2k}\frac{1}{v^{-1}\pm q^{k}}\right).$
The proposition follows. ∎
Proposition 8.2.
Let $T\colon\mathcal{A}\to\mathbb{C}$ be a $g$-twisted trace and $S\colon V\to\mathbb{C}$ be the function from Proposition 8.1. Then
$S(F(z))=\int_{\varepsilon S^{1}}F(z)\omega(z)\frac{{\rm d}z}{z},$
where $S^{1}$ is the unit circle, $\omega(z)$ is a holomorphic function on $\mathbb{C}^{\times}$ such that $\omega(qz)=q^{-\frac{m}{2}}z^{-m}\omega(z)$ and $\omega(z)=\omega\bigl(z^{-1}\bigr)$, and $\varepsilon$ is any number smaller than all of the absolute values of the poles of $F$. Conversely, every such $S$ defines a trace by the rule $T\bigl(\sum_{i}w^{i}f_{i}\bigr)=S(f_{0})$. In particular, the dimension of the space of traces $T$ that vanish on $\mathcal{A}_{1}$ is $1+\frac{m}{2}$.
@@
The proof consists of three steps. First, we rewrite the trace condition for $T|_{\mathcal{A}_{0}}$ in terms of $S$ and get equation (8.1). Then we check that each $S$ as in the statement of the proposition satisfies (8.1) and show that the space of possible functions $\omega(z)$ has dimension $\frac{m}{2}+1$. Finally, we show that the dimension of the space of $S\colon V\to\mathbb{C}$ satisfying (8.1) has dimension at most $\frac{m}{2}+1$}.
We use the trace condition $T(ba)=T(ag(b))$ for
$a=\sum w^{2i+1}f_{2i+1}\qquad\text{and}\qquad b=w\smash{\frac{v^{k}}{v-v^{-1}}}+w^{-1}\smash{\frac{v^{-k}}{v-v^{-1}}}.$
Since $ba$ and $ag(b)$ belong to $\mathcal{A}_{0}$, by Proposition 8.1, we have
$T(ba)=S\bigl(\big[w^{0}\big](ba)\bigr)\qquad\text{and}\qquad T(ag(b))=S\bigl(\big[w^{0}\big]\bigl(ag(b)\bigr)\bigr).$
We note that the only condition on $f_{1}(v)$ is that it is regular at $v=\pm q$. We have
$f_{-1}(v)=-f_{1}\bigl(v^{-1}\bigr).$
Denote $f_{1}(v)=F(v)$, so that $f_{-1}(v)=-F\bigl(v^{-1}\bigr)$. We have
$\big[w^{0}\big](ba)$ $=w\frac{v^{k}}{v-v^{-1}}\cdot\bigl(-w^{-1}F\bigl(v^{-1}\bigr)\bigr)+w^{-1}\frac{v^{-k}}{v-v^{-1}}wF(v)$
$=q^{k}\bigg(\frac{-v^{k}}{qv-q^{-1}v^{-1}}F\bigl(v^{-1}\bigr)+\frac{v^{-k}}{q^{-1}v-qv^{-1}}F(v)\bigg).$
Note that $g(b)=q^{-\frac{m}{2}}\bigl(w\frac{v^{k+m}}{v-v^{-1}}+w^{-1}\frac{v^{-k-m}}{v-v^{-1}}\bigr)$. Hence,
$\big[w^{0}\big](ag(b))$ $=q^{-\frac{m}{2}}\bigg(wF(v)w^{-1}\frac{v^{-k-m}}{v-v^{-1}}-w^{-1}F\bigl(v^{-1}\bigr)w\frac{v^{k+m}}{v-v^{-1}}\bigg)$
$=q^{-\frac{m}{2}}\bigg(F(qv)\frac{v^{-k-m}}{v-v^{-1}}-F\bigl(qv^{-1}\bigr)\frac{v^{k+m}}{v-v^{-1}}\bigg).$